# Network reconstruction via the minimum description length principle

### v1

@@ Abstract
A fundamental problem associated with the task of network reconstruction from dynamical or behavioral data consists in determining the most appropriate model complexity in a manner that prevents overfitting, and produces an inferred network with a statistically justifiable number of edges and their weight distribution. The status quo in this context is based on $L_{1}$ regularization combined with cross-validation. As we demonstrate, besides its high computational cost, this commonplace approach unnecessarily ties the promotion of sparsity, i.e. abundance of zero weights, with weight “shrinkage.” This combination forces a trade-off between the bias introduced by shrinkage and the network sparsity, which often results in substantial overfitting even after cross-validation. In this work, we propose an alternative nonparametric regularization scheme based on hierarchical Bayesian inference and weight quantization, which does not rely on weight shrinkage to promote sparsity. Our approach follows the minimum description length (MDL) principle, and uncovers the weight distribution that allows for the most compression of the data, thus avoiding overfitting without requiring cross-validation. The latter property renders our approach substantially faster and simpler to employ, as it requires a single fit to the complete data, instead of many fits for multiple data splits and choice of regularization parameter. As a result, we have a principled and efficient inference scheme that can be used with a large variety of generative models, without requiring the number of reconstructed edges and their weight distribution to be known in advance. In a series of examples, we also demonstrate that our scheme yields systematically increased accuracy in the reconstruction of both artificial and empirical networks. We highlight the use of our method with the reconstruction of interaction networks between microbial communities from large-scale abundance samples involving in the order of $10^{4}$ to $10^{5}$ species, and demonstrate how the inferred model can be used to predict the outcome of potential interventions and tipping points in the system.
@@ I Introduction
In this work we frame network reconstruction as a statistical inference problem [12], where we assume that the observed data has been sampled from a generative statistical model that has a weighted network as part of its set of parameters. In this case, the reconstruction task consists of estimating these parameters from data. This generative framing contrasts with other inferential approaches based on e.g. Granger causality [13] and transfer entropy [14] in important ways: 1. It allows for a principled framework that does not rely on ad hoc and often arbitrary confidence thresholds; 2. Can reliably distinguish between direct and indirect interactions, to the extent that this information is obtainable from the data; 3. Enables uncertainty quantification and comparison between alternative generative hypotheses; 4. Produces a generative model as a final result, which can be used not only for interpretation, but also to predict the outcomes of interventions and alternative conditions previously unseen in the data. These same features make this type of approach preferable also to non-inferential heuristics such as correlation thresholding [15, 16], known to suffer from serious deleterious artifacts [17, 12].
@@
A central obstacle to the effective implementation of inferential network reconstruction is the need for statistical regularization: A network with $N$ nodes can in principle have up to $O(N^{2})$ edges, and elementary maximum likelihood approaches do not allow us to distinguish between the non-existence of an edge and the existence of an edge with a low weight magnitude, resulting in the inference of maximally dense networks, even when the true underlying network is sparse, i.e. its number of edges is far smaller than the number of possible edges. In this case, unregularized maximum likelihood would massively overfit, incorporating a substantial amount of statistical noise it its description, and thus would not meaningfully represent the true underlying system. Becase of this, a robust inferential framework needs to include the ability to determine the most appropriate model complexity from data, which in this context means the right number of edges and their weight distribution, in a manner that does not overfit.
@@
Perhaps the most common strategy to avoid overfitting is $L_{1}$ regularization, where a penalty proportional to the sum of the weight magnitudes is subtracted from the likelihood, multipiled by a regularization parameter. This approach is a standard technique in statistics and machine learning, perhaps most well known as the central ingredient of LASSO regression [18], and has been used extensively as well for network reconstruction using multivariate Gaussian models [19, 20, 21, 22], most famously as part of the GLASSO algorithm [23], but also for the inverse Ising model [24, 25], among other problem instances.
@@
The $L_{1}$ regularization scheme has certain attractive properties: 1. It is easy to implement; 2. It results in the existence of exact zeros for the reconstructed weights; and 3. The $L_{1}$ penalty function is convex. When combined with a likelihood function that is also convex (such as the one for the multivariate Gaussian [23] or the Ising model [25]) the latter property means that this scheme yields an overall convex objective function, thus enabling the use of very efficient algorithms available only for this simpler class of optimization problems. On the other hand, this scheme also has significant limitations: 1. The regularization parameter, which determines the final sparsity of the reconstruction, needs to be provided prior to inference, requiring the desired level of sparsity to be known beforehand; 2. It causes “shrinkage” of the weights, i.e. the weight of the non-zero edges also decrease in magnitude, introducing a bias. The first disadvantage undermines this regularization approach in a central way, since it requires as an input what it should in fact provide as an output: the network sparsity. Because of this, the scheme is often used together with cross-validation [23, 26, 27] to determine the optimal value of the regularization parameter. But not only does this significantly increase the computational cost of the approach, but in the end also typically results in substantial overfitting, as we demonstrate in this work.
@@
Other regularization approaches also have been proposed for network reconstruction, such as $\delta$-thresholding [28], decimation [29], BIC [30], weight integration [31], alternative priors [32, 33, 34] with fixed number of edges [35], as well as spike-and-slab [36, 37] and horseshoe [38, 39, 40] shapes. These alternatives address some limitations of $L_{1}$, in particular its detrimental shrinkage properties, but they all either still require cross validation, rely on heuristics to determine the resulting sparsity, and/or make strong assumptions on the shape of the weight distribution.
@@ II Inferential framework
with $\bm{x}_{m}$ being the $m$-th column of $\bm{X}$. Additionally, we may also have that the network generates a Markovian time series with likelihood
@@ III $L_{1}$ regularization, shrinkage, and overfitting
Figure 1: $L_{1}$ regularization overfits when combined with cross validation. This example considers the reconstruction of the weighted Karate club network [45] ($N=34$ nodes and $E=78$ edges, weight values sampled i.i.d. from a normal distribution with mean $0.22$ and standard deviation $0.01$), based on $M=1,000$ transitions from the kinetic Ising model with a random initial state, using Eqs. 6 and 7, for a range of values of the regularization strength $\lambda$. Panel (a) shows, from top to bottom, the Jaccard similarity between inferred and true weights and binarized edges, the mean held-out likelihood for a 5-fold cross validation, the number of inferred non-zero edges, and the individual values of inferred weights (with true nonzero edges shown in blue, and true zero-valued entries shown in red). The grey horizontal lines at the right margin of the bottom of panel (a) show the true weight values. The vertical dashed lines mark the values of $\lambda$ that maximize (b) the mean held-out likelihood and (c) the binarized Jaccard similarity. The inferred network for these two values of $\lambda$ are shown in panels (b) and (c), respectively, with edge weights represented as thickness and the color representing wether it is a true (blue) or spurious (red) edge. In panel (a), all dashed horizontal lines, as well as the red line at the right margin of the bottom panel, mark the results obtained with the MDL regularization of Sec IV.
@@
It is clear from this example that the $L_{1}$ regularization obtained with the Laplace prior forces an unnecessary dichotomy between sparsity and low weight magnitudes. Ideally, we should be able to determine which edges are better modeled by setting them to zero, without inherently reducing the magnitude of the non-zero ones. This is precisely what we achieve in the next section.
@@ IV Weight quantization and the minimum description length principle
With the objective of precisely quantifying the model complexity, we proceed in several steps. First, we account for the sparsity of the model by introducing an auxiliarty variable $\bm{A}$ corresponding to the binary adjacency matrix that determines the values of $\bm{W}$ that are going to be non-zero. At first, we sample $\bm{A}$ from a uniform distribution conditioned on the total number of edges $E$,
@@
where $Y_{ij}$ is some continuous auxiliary value, and $\Delta$ represents the precision considered, such that we have a conditional probability mass $P(W_{ij}|Y_{ij},\Delta)=\delta_{W_{ij},\Delta\left\lceil Y_{ij}/\Delta\right\rfloor}$. Note that we do not assume that the true values of $W_{ij}$ are actually sampled in this way; instead this prior simply articulates the exact precision $\Delta$ with which we are performing the inference.
@@
where the remaining quantities $\bm{m}$, $\bm{z}$, $K$, $\bm{A}$, and $E$ in Eq. 27 should be interpreted as being functions of $\bm{W}$. Note that this model recovers the one using Eq. 19 when $K=E$, i.e. every weight belongs to its own category, up to a unimportant factor $1/(E+1)$—although we expect to obtain significantly smaller description length values with Eq. 27, since it can exploit regularities in the weight distribution and determine the most appropriate precision of the weights. We emphasize that, differently from Eq. 19, in Eq. 27 the parameter $\Delta$ controls the precision of the weight categories $\bm{z}$, not directly of the weight themselves. The precision of the weights $\bm{W}$ are primarily represented by the actual finite categories $\bm{z}$.
@@ V Inference algorithm
Whenever Eq. 12 corresponds to a convex optimization objective, it can be performed with relatively straighforward algorithms. The simplest one is coordinate descent [48], in which every entry in the $\bm{W}$ matrix is optimized in sequence, while keeping the others fixed. This yields an algorithm with a complexity of $O(\tau N^{2})$, provided the optimization of each entry can be done in time $O(1)$ (e.g. via bisection search), and $\tau$ is the number of iterations necessary for sufficient convergence. A recent significant improvement over this baseline was proposed in Ref. [43], where an algorithm was presented that solves the same optimization problem in subquadratic time, typically with a $O(N\log^{2}N)$ complexity, although the precise scaling will depend on the details of the data in general. This algorithm works as a greedy version of coordinate descent, where a subset of the entries are repeatedly proposed to be updated according to the NNDescent algorithm [49], which implements a stochastic search that involves keeping a list of candidate edges incident on each node in the network as an endpoint, and updating this list by investigating the list of second neighbors in the candidate network. This algorithm typically converges in log-linear time, allowing the overall reconstruction algorithm also to converge in subquadratic time, in this way converging orders of magnitude faster than the quadratic baseline of the coordinate descent algorithm.
@@
    $W_{ij}^{\prime}\in\bm{z}$ chosen according to a random bisection search maximizing the objective function, with each midpoint sampled uniformly at a random.
@@
    $W_{ij}^{\prime}\in\mathbb{R}$ and $W_{ij}^{\prime}\notin\bm{z}$ chosen according to a bisection search in the interval of weights allowed for the problem. This will introduce a new weight category with $\bm{z}^{\prime}=\bm{z}\cup\{W_{ij}^{\prime}\}$.
@@
The workhorse of this procedure is the algorithm of Ref. [43] which reduces the search for candidate updates to a subquadratic time. The remaining steps of the algorithm operate on this set with linear complexity, and therefore a single “sweep” of all latent variables can be accomplished under the same algorithmic complexity as Ref. [43], since it dominates. In practice, due to the extra latent variables and optimization steps, this algorithm finishes in time around 5 to 10 times longer than when using $L_{1}$ regularization for a single value of $\lambda$, but is much faster than doing cross-validation and scanning over many values of $\lambda$. A reference C++ implementation of this algorithm is freely available as part of the graph-tool Python library [51].
@@ V.1 Assessment of the algorithm with synthetic data
We observe that our MDL regularization performs very closely to the true prior, except for very few data, in which case the MDL tends to produce an empty network, whereas the true prior in fact overfits and yields a network with a number of edges larger than the true value. As we already explained, MAP estimation does not guarantee proper regularization when the priors correspond to probability densities—even if it happens to be the correct one. The overall tendency of the MDL approach is, by design, to err towards simpler models, instead of more complex ones, when the data is insufficient.
@@
Figure 2: Inference results for the kinetic and equilibrium Ising model for two empirical networks, as indicated in the legend, with true weights sampled as described in the text, using both the true prior and our MDL regularization. The individual panels show the Jaccard similarity between the inferred and true networks, as well as the number of inferred non-zero edges (the dashed horizontal line shows the true value). The square symbols show the MDL result obtained by initializing the algorithm with an empty network, and the triangles when initializing with the true network.
Fig. 2 also compares the result of our MDL algorithm when initialized with an empty network and with the true network. In most cases the results are identical, indicating that our algorithm does not tend to get trapped in local minima—unless in some situations with only very few data (e.g. Fig 2b), in which case it could benefit from alternative initialization strategies and/or multiple runs.
We also compare in Fig. 1 the result of the MDL regularization with $L_{1}$ for the Karate Club example, where we obtain a performance with MDL that is superior to $L_{1}$ for any value of $\lambda$.

### v3

@@ Abstract
A fundamental problem associated with the task of network reconstruction from dynamical or behavioral data consists in determining the most appropriate model complexity in a manner that prevents overfitting, and produces an inferred network with a statistically justifiable number of edges and their weight distribution. The status quo in this context is based on $L_{1}$ regularization combined with cross-validation. However, besides its high computational cost, this commonplace approach unnecessarily ties the promotion of sparsity, i.e. abundance of zero weights, with weight “shrinkage.” This combination forces a trade-off between the bias introduced by shrinkage and the network sparsity, which often results in substantial overfitting even after cross-validation. In this work, we propose an alternative nonparametric regularization scheme based on hierarchical Bayesian inference and weight quantization, which does not rely on weight shrinkage to promote sparsity. Our approach follows the minimum description length (MDL) principle, and uncovers the weight distribution that allows for the most compression of the data, thus avoiding overfitting without requiring cross-validation. The latter property renders our approach substantially faster and simpler to employ, as it requires a single fit to the complete data, instead of many fits for multiple data splits and choice of regularization parameter. As a result, we have a principled and efficient inference scheme that can be used with a large variety of generative models, without requiring the number of reconstructed edges and their weight distribution to be known in advance. In a series of examples, we also demonstrate that our scheme yields systematically increased accuracy in the reconstruction of both artificial and empirical networks. We highlight the use of our method with the reconstruction of interaction networks between microbial communities from large-scale abundance samples involving in the order of $10^{4}$ to $10^{5}$ species, and demonstrate how the inferred model can be used to predict the outcome of potential interventions and tipping points in the system.
@@ I Introduction
In this work we frame network reconstruction as a statistical inference problem [12], where we assume that the observed data has been sampled from a generative statistical model that has a weighted network as part of its set of parameters. In this case, the reconstruction task consists of estimating these parameters from data. This generative framing contrasts with other inferential approaches based on e.g. Granger causality [13] and transfer entropy [14] in important ways: 1. It allows for a principled framework that does not rely on ad hoc and often arbitrary confidence thresholds; 2. Can reliably distinguish between direct and indirect interactions, to the extent that this information is obtainable from the data; 3. Enables uncertainty quantification and comparison between alternative generative hypotheses; 4. Produces a generative model as a final result, which can be used not only for interpretation, but also to predict the outcomes of interventions and alternative conditions previously unseen in the data. These same features make this type of approach preferable also to non-inferential heuristics such as correlation thresholding [15, 16], known to suffer from substantial biases [17, 12].
@@
A central obstacle to the effective implementation of inferential network reconstruction is the need for statistical regularization: A network with $N$ nodes can in principle have up to $O(N^{2})$ edges, and elementary maximum likelihood approaches do not allow us to distinguish between the non-existence of an edge and the existence of an edge with a low weight magnitude, resulting in the inference of maximally dense networks, even when the true underlying network is sparse, i.e. its number of edges is far smaller than the number of possible edges. In this case, unregularized maximum likelihood would massively overfit, incorporating a substantial amount of statistical noise in its description, and thus would not meaningfully represent the true underlying system. Becase of this, a robust inferential framework needs to include the ability to determine the most appropriate model complexity from data, which in this context means the right number of edges and their weight distribution, in a manner that does not overfit.
@@
Perhaps the most common strategy to avoid overfitting is $L_{1}$ regularization [18], where a penalty proportional to the sum of the weight magnitudes is subtracted from the likelihood, multipiled by a regularization parameter. This approach is a standard technique in statistics and machine learning, perhaps most well known as the central ingredient of LASSO regression [19], and has been used extensively as well for network reconstruction using multivariate Gaussian models [20, 21, 22, 23], most famously as part of the GLASSO algorithm [24], but also for the inverse Ising model [25, 26, 27, 28, 29, 30, 31], among other problem instances.
@@
The $L_{1}$ regularization scheme has certain attractive properties: 1. It is easy to implement; 2. It results in the existence of exact zeros for the reconstructed weights; and 3. The $L_{1}$ penalty function is convex. When combined with a likelihood function that is also convex (such as the one for the multivariate Gaussian [24] or the Ising model [29]) the latter property means that this scheme yields an overall convex objective function, thus enabling the use of very efficient algorithms available only for this simpler class of optimization problems. On the other hand, this scheme also has well-known significant limitations [18]: 1. The regularization parameter, which determines the final sparsity of the reconstruction, needs to be provided prior to inference, requiring the desired level of sparsity to be known beforehand; 2. It causes “shrinkage” of the weights, i.e. the weights of the non-zero edges also decrease in magnitude, introducing a bias. The first disadvantage undermines this regularization approach in a central way, since it requires as an input what it should in fact provide as an output: the network sparsity. Because of this, the scheme is often used together with cross-validation [24, 32, 33] to determine the optimal value of the regularization parameter. But not only does this significantly increase the computational cost of the approach, but in the end also typically results in substantial overfitting, as we demonstrate in this work.
@@
Other regularization approaches also have been proposed for network reconstruction, such as $\delta$-thresholding [34], decimation [35, 36], EBIC [37], adaptive LASSO [38, 39, 40], MCP [41], SCAD [42], weight integration [43], alternative priors [44, 45, 46] with fixed number of edges [47], as well as spike-and-slab [48, 49] and horseshoe [50, 51, 52] shapes. These alternatives address some limitations of $L_{1}$, in particular its detrimental shrinkage properties, but they all either still require cross validation, rely on heuristics to determine the resulting sparsity, and/or make strong assumptions on the shape of the weight distribution.
@@ II Inferential framework
with $\bm{x}_{m}$ being the $m$-th column of $\bm{X}$. Alternatively, we may have that the network generates a Markovian time series with likelihood
@@ III $L_{1}$ regularization, shrinkage, and overfitting
Figure 1: $L_{1}$ regularization overfits when combined with cross validation. This example considers the reconstruction of the weighted Karate club network [57] ($N=34$ nodes and $E=78$ edges, weight values sampled i.i.d. from a normal distribution with mean $0.22$ and standard deviation $0.01$), based on $M=1,000$ transitions from the kinetic Ising model with a random initial state, using Eqs. 6 and 7, for a range of values of the regularization strength $\lambda$. Panel (a) shows, from top to bottom, the Jaccard similarity between inferred and true weights and binarized edges, the mean held-out likelihood for a 5-fold cross validation, the number of inferred non-zero edges, and the individual values of inferred weights (with true nonzero edges shown in blue, and true zero-valued entries shown in red). The grey horizontal lines at the right margin of the bottom of panel (a) show the true weight values. The vertical dashed lines mark the values of $\lambda$ that maximize (b) the mean held-out likelihood and (c) the binarized Jaccard similarity. The inferred network for these two values of $\lambda$ are shown in panels (b) and (c), respectively, with edge weights represented as thickness and the color representing whether it is a true (blue) or spurious (red) edge. In panel (a), all dashed horizontal lines, as well as the red line at the right margin of the bottom panel, mark the results obtained with the MDL regularization of Sec IV.
@@
It is clear from this example that the $L_{1}$ regularization obtained with the Laplace prior forces an unnecessary dichotomy between sparsity and low weight magnitudes. The detrimental bias of $L_{1}$ to obtain regression coefficients is well known [42] and alternative such as adaptive LASSO [38, 39, 40], MCP [41], and SCAD [42] exist which mitigate this problem; but they all involve ad hoc parameters that need to be determined via cross-validation, without the algorithmic attractiveness of $L_{1}$, since these alternatives are no longer convex. As a result, these alternatives are substantially more cumbersome to use.
Ideally, we should be able to determine which edges are better modeled by setting them to zero, without inherently reducing the magnitude of the non-zero ones. This is an inference problem on its own, and instead of committing beforehand to which kind of weight distribution is more appropriate, a more robust approach consists in formulating an open-ended class of distributions, and learning its precise shape a posteriori from the data, following the principle of maximum parsimony. This is precisely what we achieve in the next section.
@@ IV Weight quantization and the minimum description length principle
With the objective of precisely quantifying the model complexity, we proceed in several steps. First, we account for the sparsity of the model by introducing an auxiliary variable $\bm{A}$ corresponding to the binary adjacency matrix that determines the values of $\bm{W}$ that are going to be non-zero. At first, we assume $\bm{A}$ is sampled from a uniform distribution conditioned on the total number of edges $E$,
@@
[…]
@@
where the remaining quantities $\bm{m}$, $\bm{z}$, $K$, $\bm{A}$, and $E$ in Eq. 27 should be interpreted as being functions of $\bm{W}$. Note that this model recovers the one using Eq. 19 when $K=E$, i.e. every weight belongs to its own category, up to a unimportant factor $1/(E+1)$—although we expect to obtain significantly smaller description length values with Eq. 27, since it can exploit regularities in the weight distribution and determine the most appropriate precision of the weights. We emphasize that, differently from Eq. 19, in Eq. 27 the parameter $\Delta$ controls the precision of the weight categories $\bm{z}$, not directly of the weights themselves. The precision of the weights $\bm{W}$ are primarily represented by the actual finite categories $\bm{z}$.
@@ V Inference algorithm
Whenever Eq. 12 corresponds to a convex optimization objective, it can be performed with relatively straighforward algorithms. The simplest one is coordinate descent [60], in which every entry in the $\bm{W}$ matrix is optimized in sequence, while keeping the others fixed. This yields an algorithm with a complexity of $O(\tau N^{2})$, provided the optimization of each entry can be done in time $O(1)$ (e.g. via bisection search), and $\tau$ is the number of iterations necessary for sufficient convergence. A recent significant improvement over this baseline was proposed in Ref. [55], where an algorithm was presented that solves the same optimization problem in subquadratic time, typically with a $O(N\log^{2}N)$ complexity, although the precise scaling will depend on the details of the data in general. This algorithm works as a greedy version of coordinate descent, where a subset of the entries are repeatedly proposed to be updated according to the NNDescent algorithm [61]. The latter implements a stochastic search that involves keeping a list of candidate edges incident on each node in the network as an endpoint, and updating this list by investigating the list of second neighbors in the candidate network. This algorithm typically converges in log-linear time, allowing the overall reconstruction algorithm also to converge in subquadratic time, in this way converging orders of magnitude faster than the quadratic baseline of the coordinate descent algorithm.
@@
    $W_{ij}^{\prime}\in\bm{z}$ chosen according to a random bisection search  [^3] Random bisection search will eventually succeed in finding the global minimum of a multimodal objective, hence it should in general be preferred over deterministic bisection, unless the objective function is convex. maximizing the objective function, with each midpoint sampled uniformly at a random.
@@
    $W_{ij}^{\prime}\in\mathbb{R}$ and $W_{ij}^{\prime}\notin\bm{z}$ chosen according to a random bisection search in the interval of weights allowed for the problem. This will introduce a new weight category with $\bm{z}^{\prime}=\bm{z}\cup\{W_{ij}^{\prime}\}$.
@@
The workhorse of this overall procedure is the algorithm of Ref. [55] which reduces the search for candidate updates to a subquadratic time. The remaining steps of the algorithm operate on this set with linear complexity, and therefore a single “sweep” of all latent variables can be accomplished under the same algorithmic complexity as Ref. [55], since it dominates. In practice, due to the extra latent variables and optimization steps, this algorithm finishes in time around 5 to 10 times longer than when using $L_{1}$ regularization for a single value of $\lambda$, but is typically much faster than doing cross-validation and scanning over many values of $\lambda$. A reference C++ implementation of this algorithm is freely available as part of the graph-tool Python library [64].
@@ V.1 Assessment of the algorithm with synthetic data
We observe that our MDL regularization performs very closely to the true prior, except for very few data, in which case the MDL tends to produce an empty network, whereas the true prior in fact overfits and yields a network with a number of edges larger than the true value. As we already explained, MAP estimation does not guarantee proper regularization when the priors correspond to probability densities—even if it happens to be the correct one. The overall tendency of the MDL approach is, by design, to err towards simpler models, instead of more complex ones, when the data is insufficient to provide a good accuracy.
As already discussed previously in Sec. III, $L_{1}$ regularization with cross-validation performs substantially worse than our MDL approach, overfitting the data with networks that are orders of magnitude denser than they should be. Unlike using the true prior or MDL, for the examples of Fig. 2 the results obtained with $L_{1}$ show no sign of converging asymptotically to true network as the data increases with $M\to\infty$. This behavior further highlights the general unsuitability of the $L_{1}$ approach even in situations where the data is abundant. We should also mention that the results obtained with $L_{1}$ took typically one to two orders of magnitude longer to be computed, due to the need to optimize over the parameter $\lambda$ using a bisection search.
@@
Figure 2: Inference results for the kinetic and equilibrium Ising model for two empirical networks, as indicated in the legend, with true weights sampled as described in the text, using both the true prior, our MDL regularization, as wel as $L_{1}$ regularization with 5-fold cross-validation. The individual panels show the Jaccard similarity $s(\bm{W},\hat{\bm{W}})$ between the inferred and true networks, as well as the number $E$ of inferred non-zero edges (the dashed horizontal line shows the true value).